OpenStudy (anonymous):

If x/y=y/z, how does (x/y)^2=x/z? I'm not asking for the proof that it's true, i'm just wondering how looking at the first equation can lead to the second

OpenStudy (anonymous):

x/y=y/z, multiply both sides by x/y and u'll get (x/y)^2=x/z :)

OpenStudy (anonymous):

ah so it is haha, thanks :)